OpenStudy (anonymous):

write an equation of the line containing a given point and parallel to the given line. Express you answer in the given form y=mx+b (-9,4);8x=9y+2

OpenStudy (anonymous):

given equation is 8x=9y+2 => 9y = 8x - 2 so, gradient = 8/9 as the second line is parallel so gradient = 8/9 now, y= 8/9 x +c------1 4 = 8/9 (-9) + c 4 = -8 + c c = 12 plugging value to 1 y = 8/9 x + 12